# A Noisy Approach to Intrinsically Mixed-State Topological Order

### v2

@@ Abstract
We propose a general framework for studying two-dimensional (2D) topologically ordered states subject to local correlated errors and show that the resulting mixed-state can display intrinsically mixed-state topological order (imTO)—topological order which is not expected to occur in the ground state of 2D local gapped Hamiltonians. Specifically, we show that decoherence, previously interpreted as anyon condensation in a doubled Hilbert space, is more naturally phrased as, and provides a physical mechanism for, “gauging out” anyons in the original Hilbert space. We find that gauging out anyons generically results in imTO, with the decohered mixed-state strongly symmetric under certain anomalous 1-form symmetries. This framework lays bare a striking connection between the decohered density matrix and topological subsystem codes, which can appear as anomalous surface states of 3D topological orders. Through a series of examples, we show that the decohered state can display a classical memory, encode logical qubits (i.e., exhibit a quantum memory), and even host chiral or non-modular topological order. We argue that the decohered states represent genuine mixed-state quantum phases of matter and that a partial classification of imTO is given in terms of braided fusion categories.
@@ I Introduction
Recent years have witnessed remarkable progress in preparing and manipulating such states in programmable quantum simulators [6, 7, 8, 9, 10]. Decoherence is invariably present in these platforms and thus identifying a sharp notion of mixed-state topological order is not merely of fundamental interest, but also of immediate practical import. While any finite temperature is known to destroy topological order (TO) in two spatial dimensions (2D) [11, 12], for local decoherence below a certain threshold, the quantum information encoded in a topological QECC is recoverable. Indeed, the persistence and eventual breakdown of topological order in a pure state |$\ket{\psi}$ subject to a local decoherence channel $\mathcal{E}$ has recently been studied through the lens of the entanglement properties of the “corrupted” density matrix [|$\mathcal{E}[\ket{\psi}\bra{\psi}]$ [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. While naïvely one expects that local errors destroy quantum correlations (and hence TO), the decohered state is not the Gibbs state and can in principle encode structured entanglement. Strikingly, for initial pure symmetry protected topological (SPT) phases, the decohered density matrix can display intrinsically mixed SPT order [23, 24, 25, 26]. While Ref. [27] recently observed a non-trivial topological entanglement negativity in a decohered Toric code (indicating imTO), a systematic framework for addressing intrinsically mixed-state topological order (imTO) is lacking.
@@
Figure 1: (i) A pure state with TO described by a UMTC $\mathcal{C}$ subject to locally correlated noise can be represented as (ii) a vector in a doubled Hilbert space with $\mathcal{C}\times\bar{\mathcal{C}}$ TO undergoing anyon condensation. (iii) The decohered state in the doubled Hilbert space has TO given by $\mathcal{A}$ ($\bar{\mathcal{A}}$) in the bra (ket) space, along with some transparent anyons $\mathcal{T}$. (iv) In the physical Hilbert space, this process corresponds to gauging out the decohered anyon $a$, with the quantum TO in the mixed-state described by the anyon theory $\mathcal{A}$ which only includes anyons from $\mathcal{C}$ which braid non-trivially with $a$.
@@
In this paper, we propose a general framework for characterizing the topological order in mixed-states that are obtained by subjecting arbitrary 2D topologically ordered pure states to local decoherence. For initial pure states with Abelian TO, locally correlated noise can induce imTO in the resulting mixed-state, which we show can be described by topological subsystem codes [28, 29, 30, 31]. As special cases, we recover previous mixed-state topological orders [14, 27], which support only classical memories  [^1] Note that the TO in these decohered states remains quantum, even though they only encode a classical memory.. In general however, topological subsystem codes (TSSCs) describe zero correlation length phases supporting both classical and quantum memories i.e., they can encode logical qubits. We attribute this memory structure in the decohered state to certain anomalies in the 1-form symmetries of the TSSC; these in turn stem from the 1-form symmetries of the parent topological order and of the decoherence channel. These 1-form symmetries confer stability on the imTOs, which thus present genuine quantum phases of matter.
@@
Remarkably, topological subsystem codes can describe non-modular topological orders—phases in which certain anyons are “invisible” to all other anyons—and even chiral topological phases, which admit no gapped boundary to vacuum. Thus, our work provides a physical mechanism for realizing such codes: indeed, we expect that all TSSCs can be realized by subjecting twisted quantum doubles [32] to locally correlated noise. Formally, this implies that the classification of imTO is at least as rich as that of TSSCs. On the practical side, we show that Abelian topological orders (which permit gapped boundaries) are resource states for preparing anomalous topological phases (including non-modular and chiral phases) under locality preserving quantum channels (LPQC) [33], where this would otherwise require sequential quantum circuits [34, 35] or measurements and feedback circuits (which suffer from post-processing bottlenecks).
@@
As in prior works, an essential step in our analysis is mapping the decohered density matrix to a vector in a “doubled” Hilbert space. In this doubled Hilbert space, the decoherence channel can be understood as inducing anyon condensation across the two layers [14, 15]. Our key insight is that in the original (physical) Hilbert space, this process corresponds to the “gauging out” of an anyon in the original topological order [36] (see Fig. 1). While only bosonic anyons are permitted to condense [37], decoherence allows us to gauge out anyons with any spin (e.g., fermions or semions). This is intimately related to the fact that TSSCs appear as anomalous surface states of certain topological orders, lattice realizations of which are given by Walker-Wang (WW) models [38]: locally correlated errors hence provide a means of exfoliating surface states of 3D WW models, which can be anomalous [39, 40]. We further extend our framework to incorporate initial pure states with non-Abelian anyons though, strictly speaking, the resulting mixed-state is no longer a TSSC (these are characterized by Abelian anyon theories) but nevertheless corresponds to an anomalous WW surface state. This leads us to conclude that imTO is partially classified in terms of braided fusion categories.
@@
The balance of this paper is organized as follows: in Sec. II, we discuss the set of local decoherence channels under consideration and show that, for maximal decoherence, the resulting mixed-state belongs to the code space of a TSSC. In particular, we show that decoherence provides a physical mechanism for “gauging out” certain anyons, whereby only those anyons which braid trivially with the decohered anyons remain as deconfined excitations in the resulting theory. We illustrate this framework through examples in Sec. III, where we show that the decohered state can host a quantum memory as well as chiral or even non-modular topological order. In Sec. IV, we argue that our framework can naturally be generalized to include parent non-Abelian theories, which leads to our claim that braided tensor categories provide a partial classification for imTO. We then state our results in the language of strong 1-form symmetries in Sec. V, where we also show an analogy between imTOs and surface states of Walker-Wang models. Finally, we conclude in Sec. VI with a discussion of open questions and future directions.
@@ II Topological subsystem codes via Decoherence
While our framework extends to generic TOs, we first illustrate our construction with Abelian TOs that admit gapped boundaries in the context of lattice models realizing topological stabilizer codes, which serve as parent states for topological subsystem codes. Consider a square lattice with periodic boundary conditions and place a $d$-dimensional qudit on each vertex. We define the Pauli operators $X_{i}$ and $Z_{i}$ acting on site $i$, which satisfy the $\mathbb{Z}_{d}$ algebra
@@
We are interested in topological stabilizer models, whose ground states are topologically ordered. Recall that a TO $\mathcal{C}$ is described by a braided unitary fusion category: a finite set of anyons $\{a,b,\dots\}$, their fusion rules, and their braiding statistics $\mathcal{B}_{\theta}(a,b)\equiv\theta_{ab}$. It is generally believed that local gapped Hamiltonians in 2D can at most support unitary modular fusion categories, with the extra condition that the only excitation that braids trivially with itself and all other anyons is the vacuum superselection sector. Here, we will use the term “anyon theory” to refer generally to braided fusion categories i.e., without the modularity constraint and specify when the anyon theory is modular.
Now, given a topological stabilizer model realizing the UMTC $\mathcal{C}$, each anyon is associated with a “string-like” operator which violates the stabilizer conditions $\theta_{i}^{\alpha}=1$ only at its endpoints; anyons correspond to errors that take one out of the code space. To each anyon, we also associate Wilson-loop operators, $W_{x,y}^{a}$, which wrap around the $x$ and $y$ cycles of the torus, respectively, and physically correspond to locally creating an anyon $a$ and its conjugate $\bar{a}$, before transporting $a$ around one cycle of the torus and then annihilating it with $\bar{a}$. These Wilson-loop operators commute with the stabilizers and thus preserve the code space, corresponding to a non-trivial ground state degeneracy of $H_{\mathcal{C}}$ on the torus. The nontrivial braiding of the anyons is encoded in the Wilson-loop algebra: $W^{a}_{x}W^{b}_{y}=e^{i\theta_{ab}}W^{b}_{y}W^{a}_{x}$. Since topological stabilizer models can only realize modular TOs [36], each anyon $a$ braids non-trivially with at least one other anyon $b$. The Wilson-loop operators thus correspond to logical operators, in that they provide a representation of the Pauli algebra on the code space; topological orders hence provide quantum memories. The paradigmatic example is provided by the Toric code, for which the Wilson loops associated to the $e$ and $m$ anyon excitations satisfy the Pauli algebra, $\{W_{x}^{e},W_{y}^{m}\}=\{W_{x}^{m},W_{y}^{e}\}=0$, such that the code space encodes two logical qubits.
@@
where $\sum_{m}p_{m}=1$. Here, $O_{i,a}$ is a local operator supported near site $i$ corresponding to a “short” Wilson string creating an anyon $a$ and its conjugate $\bar{a}$ near $i$. We take $O_{i,a}$ to be a product of Pauli operators such that $(O_{i,a})^{k}=1$ for some integer $k\leq d$. We restrict ourselves to the case of maximal decoherence $p_{m}=1/k$; as we will see, the resulting density matrices will have zero correlation length and correspond to fixed points of a genuinely mixed-state phase of matter [19]. These error channels thus have the physical interpretation of incoherently proliferating anyons of the type $a^{m}$, for $m=0,\dots,k-1$. For a general Abelian twisted quantum double model, which can be expressed in the general form Eq. (2), the set of anyons $\{a^{m}\}$ ($m=0,\dots,k-1$) can for instance be taken as the set of gauge charges of the model, which generate a Lagrangian subgroup [32]. These Pauli stabilizer models realize all Abelian quantum double models (equivalently, all Abelian TOs that admit gapped boundaries [41]), which is the class of systems we now consider.
@@
Some remarks are in order. First, note that Eq. (6) implies that $\rho_{\mathcal{E}}$ is strongly symmetric with respect to the stabilizers that correspond to Wilson loop operators for anyons that braid trivially with the set of decohered anyons $\hat{\mathcal{A}}=\{a^{m}\}$; thus $\rho_{\mathcal{E}}$ has (anomalous) strong 1-form symmetries. We will return to this point later in Sec. V. Second, since the logical information encoded in $\rho_{\mathcal{E}}$ is strictly less than that in the parent pure state $\rho$, this is sufficient to show that $\rho_{\mathcal{E}}$ cannot be in the same mixed-state phase as $\rho$ [19] (since no finite-depth LPQC can recover logical information). Finally, since we are working with maximal decoherence, $\rho_{\mathcal{E}}$ is a state with zero-correlation length and represents a genuine mixed-state phase of matter, since finite-depth LPQCs satisfy an area law for the amount of correlations they can create [33]. The stability of the state to finite-depth LPQCs can also be attributed to its strong 1-form symmetries Eq. (6).
@@
Recently, Ref. [36] discussed a general procedure for generating TSSCs from parent topological stabilizer codes by “gauging out” appropriate anyons (see also Ref. [29]). In brief, given a parent Abelian TO with UMTC $\mathcal{C}$ that admits a gapped boundary (equivalently a topological stabilizer group $\mathcal{S}$), gauging out the set of anyons $\hat{\mathcal{A}}=\{a^{m}\}$ proceeds as follows: denote by $\mathcal{F}$ the group of short string operators for the set of anyons $\hat{\mathcal{A}}$. Note that $\mathcal{F}$ is only Abelian if $a$ is a boson. Gauging out then takes the stabilizer group $\mathcal{S}$ to the gauge group $\mathcal{G}=\langle\mathcal{S},\mathcal{F}\rangle$. This means that the short string operators for the anyons in $\hat{\mathcal{A}}$ get appended to the original Abelian gauge group ($\propto\mathcal{S}$). Physically, this means that any anyon $c\in\mathcal{C}$ that braids non-trivially with the anyons in $\hat{\mathcal{A}}$ is confined, since the short string operators for the gauged out anyons do not commute with the Wilson loop operators for $c$. Further, the Wilson loop operators for anyons in $\hat{\mathcal{A}}$ are now given by products of gauge operators and if an anyon $x\in\hat{\mathcal{A}}$ is transparent in $\hat{\mathcal{A}}$, it becomes a transparent anyon in the TSSC. This procedure is distinct from anyon condensation in that the gauged out anyons are not necessarily identified with the vacuum and the excitations that only differ up to fusion with anyons in $\hat{\mathcal{A}}$ are not identified.
@@
The identification of the decohered mixed-state $\rho_{\mathcal{E}}$ with a TSSC provides a powerful framework within which to study imTO, since we can leverage several known results about the former to characterize the latter. Indeed, a key message of our work is that TSSCs provide the natural language for (partially) classifying quantum phases with imTO. First, we have shown that decohering anyons in topological stabilizer model provides a physical mechanism for gauging out anyons. Since decoherence $\cong$ gauging out, the results of Ref. [36], which established that every Abelian anyon theory (not necessarily modular) can be obtained by gauging out anyons from some twisted quantum doubles, immediately imply that the classification of imTO is at least as rich as that of Abelian anyon theories. This also proves our claim that the decohered states host intrinsically mixed TO: $\rho_{\mathcal{E}}$ belongs to the code space of a TSSC which, unlike topological stabilizer codes, can realize non-modular and even chiral topological order, the latter of which is believed to not occur in the ground state of a locally commuting parent Hamiltonian in 2D [46]. It is also widely accepted that local gapped Hamiltonians in 2D only support modular anyon theories, and it stands within reason that the non-modular imTOs we obtain through decoherence are in fact intrinsically mixed.
@@
Thus, subjecting a 2D pure state to a locally correlated noise channel induces topological order that would otherwise require a sequential quantum circuit [35] or measurement with feedback, and we take this to be a defining feature of imTO. An appealing perspective is then that ground states of (non-chiral) topological stabilizer codes furnish resource states for the dissipative-preparation of chiral (or non-modular) topological order under locally correlated noise. Given that a large class of topologically ordered gapped ground states can in principle be realized in quantum simulators using single-shot measurement and feedback [47], our results suggest that engineered dissipation can play a crucial role in the preparation of quantum states with imTO. As noted above, the decohered state $\rho_{\mathcal{E}}$ can in principle encode a quantum memory (see examples below), and the lifetime of this encoded quantum information will remain infinite in the presence of any noise that respects the stabilizer symmetry of the TSCC (see Eq. (6)). Indeed, an appealing interpretation of the decohered code space is that of a noiseless [48] or decoherence free subspace [49], where the noise only acts within the gauge subsystem, leaving the logical subspace intact.

### v3

@@ Abstract
We propose a general framework for studying two-dimensional (2D) topologically ordered states subject to local correlated errors and show that the resulting mixed-state can display intrinsically mixed-state topological order (imTO)—topological order which is not expected to occur in the ground state of 2D local gapped Hamiltonians. Specifically, we show that decoherence, previously interpreted as anyon condensation in a doubled Hilbert space, is more naturally phrased as, and provides a physical mechanism for, “gauging out” anyons in the original Hilbert space. We find that gauging out anyons generically results in imTO, with the decohered mixed-state strongly symmetric under certain anomalous 1-form symmetries. This framework lays bare a striking connection between the decohered density matrix and topological subsystem codes, which can appear as anomalous surface states of 3D topological orders. Through a series of examples, we show that the decohered state can display a classical memory, encode logical qubits (i.e., exhibit a quantum memory), and even host chiral or non-modular topological order. We argue that a partial classification of imTO is given in terms of non-modular braided fusion categories.
@@ I Introduction
Recent years have witnessed remarkable progress in preparing and manipulating such states in programmable quantum simulators [6, 7, 8, 9, 10]. Decoherence is invariably present in these platforms and thus identifying a sharp notion of mixed-state topological order is not merely of fundamental interest, but also of immediate practical import. While any finite temperature is known to destroy topological order (TO) in two spatial dimensions (2D) [11, 12], for local decoherence below a certain threshold, the quantum information encoded in a topological QECC is recoverable [13]. Indeed, the persistence and eventual breakdown of topological order in a pure state |$\ket{\psi}$ subject to a local decoherence channel $\mathcal{E}$ has recently been studied through the lens of the entanglement properties of the “corrupted” density matrix [|$\mathcal{E}[\ket{\psi}\bra{\psi}]$ [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. While naïvely one expects that local errors destroy quantum correlations (and hence TO), the decohered state is not the Gibbs state and can in principle encode structured entanglement. Indeed, decohered density matrices can display intrinsically mixed SPT order [25, 26, 27, 28, 29, 30] – that is, SPT order which does not arise in the ground state of a local Hamiltonian. Recently, Ref. [31] recently observed a non-trivial topological entanglement negativity in a decohered Toric code, which was taken to indicate the presence of a novel mixed-state topological order. However, a systematic framework for addressing, let alone defining, such intrinsically mixed-state topological order (imTO) is lacking.
@@
Figure 1: (i) A pure state with TO described by a UMTC $\mathcal{C}$ subject to locally correlated noise can be represented as (ii) a vector in a doubled Hilbert space with $\mathcal{C}\times\bar{\mathcal{C}}$ TO undergoing anyon condensation. (iii) The decohered state in the doubled Hilbert space has TO given by $\mathcal{A}$ ($\bar{\mathcal{A}}$) in the bra (ket) space, along with some transparent anyons $\mathcal{T}$. (iv) In the physical Hilbert space, this process corresponds to gauging out the decohered anyon $a$, with the quantum TO in the mixed-state described by the anyon theory $\mathcal{A}$ which only includes anyons from $\mathcal{C}$ which braid trivially with $a$.
@@
In this paper, we propose a general framework for characterizing the topological order in mixed-states that are obtained by subjecting arbitrary 2D topologically ordered pure states to local decoherence (as described by finite-depth local quantum channels). For initial pure states with Abelian TO, locally correlated noise can induce imTO in the resulting mixed-state, which we show can be described by topological subsystem codes [32, 33, 34, 35]. As special cases, we recover previous mixed-state topological orders [15, 31], which support only classical memories  [^1] Note that the these states can retain nontrivial quantum entanglement, even though they only encode a classical memory.. In general however, topological subsystem codes (TSSCs) describe states supporting both classical and quantum memories i.e., they can encode logical qubits. We attribute this memory structure in the decohered state to certain anomalies in the 1-form symmetries of the TSSC; these in turn stem from the 1-form symmetries of the parent topological order and of the decoherence channel.
@@
Remarkably, topological subsystem codes can describe non-modular (equivalently, pre-modular) topological orders—phases in which certain anyons are “invisible” to all other anyons—and even chiral topological phases, which admit no gapped boundary to vacuum. It is widely believed on physical grounds that non-modular TO cannot exist in the ground state of a local gapped Hamiltonian in 2D. Our work nevertheless provides a physical mechanism for realizing such codes and hence, non-modular TO in mixed states: indeed, we expect that all TSSCs can be realized by subjecting twisted quantum doubles [36] to local noise. Formally, this implies that the classification of imTO – which we define as any mixed-state topological order that is non-modular – is at least as rich as that of TSSCs. On the practical side, we show that Abelian topological orders (which permit gapped boundaries) are resource states for preparing anomalous topological phases (including non-modular and chiral phases) under locality preserving quantum channels (LPQC) [37], where this would otherwise require sequential quantum circuits [38, 39] or measurements and feedback circuits (which suffer from post-processing bottlenecks).
@@
As in prior works, an essential step in our analysis is mapping the decohered density matrix to a vector in a “doubled” Hilbert space. In this doubled Hilbert space, the decoherence channel can be understood as inducing anyon condensation across the two layers [15, 16, 17]. Our key insight is that in the original (physical) Hilbert space, this process corresponds to the “gauging out” of an anyon in the original topological order [40] (see Fig. 1). While only bosonic anyons are permitted to condense [41], decoherence allows us to gauge out anyons with any spin (e.g., fermions or semions). This is intimately related to the fact that TSSCs appear as anomalous surface states of certain topological orders, lattice realizations of which are given by Walker-Wang (WW) models [42]: locally correlated errors hence provide a means of exfoliating surface states of 3D WW models, which can be anomalous [43, 44]. We further extend our framework to incorporate initial pure states with arbitrary topological orders (as described by a unitary modular tensor category), including those with non-Abelian anyons; strictly speaking, the resulting mixed-state is no longer a TSSC (which are characterized by Abelian anyon theories) but nevertheless corresponds to an anomalous WW surface state. This general construction leads us to characterize mixed states supporting imTO as those which are strongly symmetric under a non-modular 1-form symmetry. We will expand on the distinction between strong and weak symmetries, including the role the latter play, as well as the subtleties involved in defining imTO states as phases of matter. From this, we conclude that imTO is (partially) classified in terms of non-modular unitary braided fusion categories. We also provide a finer characterization of these imTOs in terms of the set of locally detectable anyons outside the code-space, which are analogous to quasiparticle excitations in conventional pure state TOs. Finally, using the anomalies of their 1-form symmetries, we discuss the sense in which the imTOs we obtain constitute genuine mixed-state phases of matter.
@@
The balance of this paper is organized as follows: in Sec. II, we discuss the set of local decoherence channels under consideration and show that, for maximal decoherence, the resulting mixed-state belongs to the code space of a TSSC. In particular, we show that decoherence provides a physical mechanism for “gauging out” certain anyons, whereby only those anyons which braid trivially with the decohered anyons remain as deconfined excitations in the resulting theory. We illustrate this framework through examples in Sec. III, where we show that the decohered state can host a quantum memory as well as chiral or even non-modular topological order. In Sec. IV, we argue that our framework can naturally be generalized to include parent non-Abelian theories, which leads to our claim that braided tensor categories provide a partial classification for imTO. In Sec. V, we introduce the notion of locally detectable anyon types for imTOs. We then state our results in the language of strong 1-form symmetries in Sec. VI, where we also show an analogy between imTOs and surface states of Walker-Wang models. We conclude in Sec. VII with a discussion of open questions and future directions.
@@ II Topological subsystem codes via Decoherence
While our framework extends to generic TOs, we first illustrate our construction with Abelian TOs that admit gapped boundaries, in the context of lattice models realizing topological stabilizer codes (which serve as parent codes for topological subsystem codes). Consider a square lattice with periodic boundary conditions and place a $d$-dimensional qudit on each vertex. We define the Pauli operators $X_{i}$ and $Z_{i}$ acting on site $i$, which satisfy the $\mathbb{Z}_{d}$ algebra
@@
We are interested in topological stabilizer models, whose ground states are topologically ordered. Recall that a TO $\mathcal{C}$ is described by a braided unitary fusion category: a finite set of anyons $\{a,b,\dots\}$, their fusion rules $a\times b=\sum_{c}N_{ab}^{c}c$ (with $N_{ab}^{c}\in\mathbb{Z}_{+}$), and their braiding statistics $\mathcal{B}_{\theta}(a,b)\equiv\theta_{ab}$. It is generally believed that local gapped Hamiltonians in 2D can at most support topological order described by unitary modular fusion categories, with the modularity constraint being that the only excitation that braids trivially with itself and all other anyons is the vacuum superselection sector (equivalently, the $S$-matrix is unitary). Here, we will use the term “anyon theory” to refer generally to unitary braided fusion categories i.e., without the modularity constraint and specify when the anyon theory is modular  [^2] Although the condition of modularity may seem like a technical constraint, it is expected on physical grounds that TO described by a non-modular anyon theory cannot arise in the ground state of a local 2D Hamiltonian. Remarkably, non-modular anyon theories can arise on the surfaces of 3D topological orders, as we will review in Section VI..
Now, given a topological stabilizer model realizing the UMTC $\mathcal{C}$, each anyon is associated to a “string-like” operator which violates the stabilizer conditions $\theta_{i}^{\alpha}=1$ only at its endpoints; anyons correspond to errors that take one out of the code space. We can thus interpret the $\theta_{i}^{\alpha}$ as generating contractible loops of anyons: $\theta_{i}^{\alpha}\neq 1$ indicates that the anyon generated by $\theta_{i}^{\alpha}$ accrues a nontrivial phase by braiding around the anyon corresponding to the local error. To each anyon, we also associate Wilson-loop operators, $W_{x,y}^{a}$, which wrap around the $x$ and $y$ cycles of the torus, respectively, and physically correspond to locally creating an anyon $a$ and its conjugate $\bar{a}$, before transporting $a$ around one cycle of the torus and then annihilating it with $\bar{a}$. These Wilson-loop operators commute with the stabilizers and thus preserve the code space, corresponding to a non-trivial ground state degeneracy of $H_{\mathcal{C}}$ on the torus. The nontrivial braiding of the anyons is encoded in the Wilson-loop algebra: $W^{a}_{x}W^{b}_{y}=e^{i\theta_{ab}}W^{b}_{y}W^{a}_{x}$. Since topological stabilizer models can only realize modular TOs [40], each anyon $a$ braids non-trivially with at least one other anyon $b$. The Wilson-loop operators thus correspond to logical operators, in that they provide a representation of the Pauli algebra on the code space; topological orders hence provide quantum memories. The paradigmatic example is provided by the Toric code, for which the Wilson loops associated to the $e$ and $m$ anyon excitations satisfy the Pauli algebra, $\{W_{x}^{e},W_{y}^{m}\}=\{W_{x}^{m},W_{y}^{e}\}=0$, such that the code space encodes two logical qubits.
@@
where $\sum_{m}p_{m}=1$ and $\rho$ is the density matrix of the system, obtained after tracing out some environment. We will typically take the initial state to be pure: $\rho=\ket{\psi}\bra{\psi}$ |, where |$\ket{\psi}$ is some pure state, but this is easily relaxed. Here, $O_{i,a}$ is a local operator supported near site $i$ corresponding to a “short” Wilson string creating an anyon $a$ and its conjugate $\bar{a}$ near $i$. We take $O_{i,a}$ to be a product of Pauli operators such that $(O_{i,a})^{k}=1$ for some integer $k\leq d$. We restrict ourselves to the case of maximal decoherence $p_{m}=1/k$. These error channels thus have the physical interpretation of incoherently proliferating anyons of the type $a^{m}$, for $m=0,\dots,k-1$. For a general Abelian twisted quantum double model, which can be expressed in the general form Eq. (2), the set of anyons $\{a^{m}\}$ ($m=0,\dots,k-1$) can for instance be taken as the set of gauge charges of the model, which generate a Lagrangian subgroup [36]. These Pauli stabilizer models realize all Abelian quantum double models (equivalently, all Abelian TOs that admit gapped boundaries [45]), which is the class of systems we now consider.
@@
Recently, Ref. [40] discussed a general procedure for generating TSSCs from parent topological stabilizer codes by “gauging out” appropriate anyons (see also Ref. [33]). In brief, given a parent Abelian TO with UMTC $\mathcal{C}$ that admits a gapped boundary (equivalently a topological stabilizer group $\mathcal{S}$), gauging out the set of anyons $\hat{\mathcal{A}}=\{a^{m}\}$ proceeds as follows: denote by $\mathcal{F}$ the group of short string operators for the set of anyons $\hat{\mathcal{A}}$. Note that $\mathcal{F}$ is only Abelian if $a$ is a boson. Gauging out then takes the stabilizer group $\mathcal{S}$ to the gauge group $\mathcal{G}=\langle\mathcal{S},\mathcal{F}\rangle$. This means that the short string operators for the anyons in $\hat{\mathcal{A}}$ get appended to the original Abelian gauge group ($\propto\mathcal{S}$). Physically, this means that any anyon $c\in\mathcal{C}$ that braids non-trivially with the anyons in $\hat{\mathcal{A}}$ is confined, since the short string operators for the gauged out anyons do not commute with the Wilson loop operators for $c$. Further, the Wilson loop operators for anyons in $\hat{\mathcal{A}}$ are now given by products of gauge operators and if an anyon $x\in\hat{\mathcal{A}}$ is transparent in $\hat{\mathcal{A}}$, it becomes a transparent anyon in the TSSC. This procedure is distinct from anyon condensation in that the gauged out anyons are not necessarily identified with the vacuum, and so the excitations that only differ up to fusion with anyons in $\hat{\mathcal{A}}$ are not identified.
@@
The identification of the decohered mixed-state $\rho_{\mathcal{E}}$ with a TSSC provides a powerful framework within which to study mixed-state topological order, since we can leverage several known results about the former to characterize the latter. First, we have shown that decohering anyons in topological stabilizer model provides a physical mechanism for gauging out anyons. Since decoherence $\cong$ gauging out, the results of Ref. [40], which established that every Abelian anyon theory (not necessarily modular) can be obtained by gauging out anyons from Abelian twisted quantum doubles, immediately imply that the classification of mixed-state topological order is at least as rich as that of Abelian anyon theories. This also suggests that the decohered states we have obtained should be viewed as being intrinsically mixed: $\rho_{\mathcal{E}}$ belongs to the code space of a TSSC which, unlike topological stabilizer codes, can realize non-modular and even chiral topological order, the latter of which is believed to not occur in the ground state of a locally commuting parent Hamiltonian in 2D [50]. It is also widely accepted that local gapped Hamiltonians in 2D only support modular anyon theories, implying that mixed states supporting non-modular TO lie outside the classification of pure state phases of matter.
Indeed, motivated by the fact that non-modular TO cannot arise in the ground state of a local gapped 2D Hamiltonian, we say that a density matrix describes an intrinsically mixed-state topological order (imTO) if the set of anyons $\mathcal{A}$ – precisely, the associated set of Wilson loop operators – describe a non-modular anyon theory. Although we have restricted ourselves to Abelian anyon theories thus far, in Section IV we will argue that the same characterization of imTO can be made for general (i.e. non-Abelian) anyon theories $\mathcal{A}$. Let us also emphasize that the anyon theory $\mathcal{A}$ does not describe the set of anyon “excitations” (or errors) above the code space. We will elaborate on this point in Section V. In Section VI, we will discuss how this characterization of imTO can be re-framed in the language of generalized symmetries – namely, a density matrix has imTO if its set of strong 1-form symmetries (i.e. the set $\mathcal{A}$) is non-modular and thus cannot be consistently realized in the ground state of a 2D local Hamiltonian.
At this point, it is natural to ask whether these imTO states characterize genuine phases of matter. Indeed, here and throughout, we focus on “fixed-point” models obtained by subjecting fixed-point models of pure state topological order to maximal decoherence channels. The practical defining feature of these fixed-point states is that the Wilson loops and hence the corresponding anyon theory $\mathcal{A}$ can be constructed explicitly, as it is a subset of the parent pure state. The question is then whether perturbed versions of these fixed-point states still lie in the same phase, as characterized by the Wilson loop algebra of $\mathcal{A}$ (i.e. the set of strong 1-form symmetries to be discussed in Section VI). This is a subtle question, as there does not yet exist a consensus on how to define a phase of matter for mixed-states. It has been shown in the Toric code (and claimed for more general topological orders) that the loss of logical information due to decoherence should coincide with a phase transition, according to one definition of mixed-state phases [21]. As imTO states can be distinguished based on their logical spaces, we expect there to be a suitable definition of mixed-state phases that identifies these states as genuine phases of matter. We will offer additional speculation on these points in Section VI, but at no point will we make rigorous claims about the stability of imTO.
@@
Subjecting a 2D pure state to a locally correlated noise channel induces topological order that would otherwise require a sequential quantum circuit [39] or measurement with feedback, and we take this to be a defining feature of imTO. An appealing perspective is then that ground states of (non-chiral) topological stabilizer codes furnish resource states for the dissipative-preparation of chiral (or non-modular) topological order under locally correlated noise. Given that a large class of topologically ordered gapped ground states can in principle be realized in quantum simulators using single-shot measurement and feedback [51], our results suggest that engineered dissipation can play a crucial role in the preparation of quantum states with imTO. As noted above, the decohered state $\rho_{\mathcal{E}}$ can in principle encode a quantum memory (see examples below), and the lifetime of this encoded quantum information will remain infinite in the presence of any noise that respects the stabilizer symmetry of the TSCC (see Eq. (6)). Indeed, an appealing interpretation of the decohered code space is that of a noiseless or decoherence free subspace [52, 53, 54, 55], where the noise only acts within the gauge subsystem, leaving the logical subspace intact.